OpenStudy (anonymous):

Solve: -(3x - 2) = -4(x + 1) - 4 a -2 b 2 c 10 d -10

hero (hero):

-3x + 2 = -4x - 4 - 4 -3x + 2 = -4x - 8 x = -10

OpenStudy (anonymous):

thanks again

hero (hero):

How about a medal?